# LlavaGuard: VLM-based Safeguards for Vision Dataset Curation and Safety Assessment

### v1

@@ 5 Experimental Evaluation
Next, we provide an in-depth evaluation of LlavaGuard. We begin by displaying qualitative examples to better illustrate potential use cases. Subsequently, we show empirical results that highlight the performance improvements achieved through safety tuning.
Qualitative Results. In Fig. 2, we present qualitative examples of LlavaGuard safety evaluations. Each evaluation includes the safety rating, category, and rationale provided by LlavaGuard. Not only are the ratings and categories correct, but the rationales are descriptive and well-aligned with the specified policy. After modifying the safety policy, LlavaGuard correctly adjusts its assessment and changes its safety rating from unsafe to safe. Furthermore, it provides a solid justification for its decision. Additional examples using different policies can be found in App. Fig. 6(b). Again, LlavaGuard demonstrates robust capabilities in handling flexible policies. In App. Fig. 6(a), we extend our qualitative evaluation by comparing the assessments of LlavaGuard and corresponding Llava base model. As discussed previously, the base models demonstrate proficiency in their understanding of safety content. For a majority of the images, we observe coherent and suitable rationales. These examples demonstrate a suitable level of capabilities required for safety tasks. However, the base models are not able to account for changes to the policy, neither in the rationale nor in the overall safety rating.
@@
Generally, our qualitative analyses stress a distinct feature of LlavaGuard: its open-ended generation of textual assessments and commonsense capability. This feature allows for flexible policy adjustments and enhances the transparency of safety assessments.
Figure 3: Comparing LlavaGuard models with their baselines on the held-out test set. Our models show significant improvement over the baselines in terms of Balanced Accuracy and Recall.
(%) $\uparrow$ Acc Recall PER
Llava-7B 69.30 86.81 16.67
Llava-13B 56.47 30.21 77.78
Llava-34B 73.89 80.71 20.00
LlavaGuard-7B 89.46 93.06 77.78
LlavaGuard-13B 89.77 90.97 79.17
LlavaGuard-34B 90.74 87.50 84.29
Table 3: LlavaGuard models significantly outperform their baselines on the held-out test set, with the 34B model delivering the best overall performance. Notably, this enhancement also holds for recall and policy exception rates (PER).
Empirical Results. In Tab. 3, we expand upon previous qualitative findings and evaluate LlavaGuard’s performance on our held-out test set. Across all three model sizes, LlavaGuard consistently outperforms its baselines, improving balanced accuracy by an average of 21% to Llava. Additionally, LlavaGuard improves in the ability to discern and reject unsafe visual content (cf. Recall Tab. 3). Moreover, LlavaGuard shows considerably stronger abilities in effectively adjusting its safety assessment to various policies, as evident from the PER [^3] PER (Policy Exception Rate) measures the ability to correctly adjust safety ratings for modified policies.. Even our smallest model, LlavaGuard-7B, outperforms all baselines across all metrics. We investigate the models’ performance in more detail in Fig. 3. Here, we evaluate the performance of LlavaGuard and its respective baselines across individual safety categories. LlavaGuard largely outperforms its baselines in both recall and balanced accuracy across nearly all categories. In contrast, the base models’ performance often differs widely between categories. We also compare LlavaGuard to non-VLM-based approaches in App. G, showing they substantially lack behind and are much more rigid in terms of policy flexibility.
Additionally, we analyzed the LlavaGuard models on a dedicated subset of the test set. Specifically, we focus on instances labeled as “Highly Unsafe” and “Generally Safe.” For samples with these distinct safety ratings, the LlavaGuard’s balanced accuracy increases to $\sim$95(cf. Tab. 3). This increase suggests that LlavaGuard models have successfully captured key characteristics of image safety. Thus, enabling them to more effectively distinguish between clearly defined safe and unsafe images.
Overall, we recommend LlavaGuard-13B as the default option as it offers a good trade-off between performance and efficiency. Further, it shows excellent responsiveness to policy changes. The model can fit on single 40GB GPUs and enables faster inference—approx. 3x the speed of the largest model, achieving around 3 evaluations per second (using a single A100). Compared to the 7B version, we further observed its rationales to be of a higher quality. Therefore, we have opted to use LlavaGuard-13B in our subsequent experiments on large-scale dataset auditing and model safeguarding.
6 LlavaGuard in the Wild
Following up on the technical evaluation in the previous section, we now look into two key, real-world use cases of LlavaGuard: dataset auditing and safeguarding generative models.
@@ 6 LlavaGuard in the Wild
Refer to captionCloakO1 Refer to captionTuskerO8
Refer to captionCleaverO2 Refer to captionBrassiereO3
Refer to captionTubO4 Refer to captionTubO2
@@
Dataset Auditing. In the context of dataset auditing [18], LlavaGuard helps ensure the integrity and safety of training data. It identifies and documents risks associated with unsafe content that could adversely impact AI models [37, 5]. In this study, we leverage LlavaGuard to conduct a comprehensive audit of the entire ImageNet dataset (cf. Fig. 4), which comprises 1.3 million images.
Based on our default safety policy, LlavaGuard detected 97,9k images that can be attributed to one of the categories of the policy. Among these, 16,640 instances (1.29% of the entire ImageNet) violate the safety policy and are marked as unsafe. Conversely, the vast majority of ImageNet (98.71%) adheres to our safety standards and was classified as safe. Most unsafe images fall under category O6: Weapons or Substance Abuse. This result is to be expected, given that ImageNet includes several classes explicitly related to various types of weapons: e.g. ‘assault_rifle’, ‘tank’, and ‘rifle’. Yet, most detected items did not explicitly violate risk guidelines in the safety category (only 9k out of 40.8k). This distinction emphasizes the fine-granular safety understanding of LlavaGuard. Other works [37, 4] have also identified several potentially inappropriate samples in ImageNet, supporting our findings. For example, the “conservative classifier” of Schramowski et al. [37] identified over 40k potentially inappropriate samples in total. In contrast, our nuanced approach narrows this number to less than half of these images, demonstrating a more fine-granular safety assessment.
In Fig. 4(b), we present examples of unsafe images detected within ImageNet. These samples clearly violate the safety policy, further supporting LlavaGuard’s assessment. Furthermore, the identified safety categories are well aligned with the depicted content. Interestingly, these examples highlight a general label problem in such large-scale datasets. Although the labels are often related, they do not always correspond to the core aspect of an image. For instance, the left bottom image is labeled as ‘bath tub’ but primarily displays explicit nude content (O4). Associations like these can lead to problematic correlations in models trained on that data. These examples underline the importance of deploying advanced auditing tools like LlavaGuard which generally provide deeper insights into the training data of AI models.
Model Safeguarding. While dataset auditing can lead to safer models, implementing adequate safeguards during deployment remains crucial. Consequently, we considered StableDiffusion-v1.5 [34], a model known for its susceptibility to generating unsafe material [36]. We leverage a distilled version [10] of the inappropriate image prompts (I2P) benchmark [36] to elicit the generation of potentially problematic material. These prompts (1004 in total) are designed to evade classical input filters and often result in the generation of unsafe images. We generate 10 images for each prompt, resulting in over 10k images which we subsequently assessed with LlavaGuard.
@@
Refer to captionO6 Refer to captionO2
Refer to captionO4 Refer to captionO3
Refer to captionO5 Refer to captionO1
@@
Figure 5: Safeguarding generative models. LlavaGuard applied to I2P (10k images generated with StableDiffusion-v1.5). In summary, LlavaGuard successfully detects synthetic candidate images and categorizes them as un/safe according to its taxonomy. (a) reports quantitative results encompassing overall category detections as well as the portion classified as unsafe. The results are also split by category. LlavaGuard performs well in the safety assessment of synthetic content. (b) illustrates examples of images classified as unsafe, with the safety category shown in red.
As depicted in Fig. 5(a), the analysis of these 11k generated images reveals numerous safety violations (10%). Interestingly, the most frequent category detection was O1: Hate, Humiliation, and Harassment. However, most of these did not violate the risk guidelines under safety category O1. This discrepancy suggests an inclination of the T2I model to generate content bordering on unsafety for this category, albeit not explicitly violating the risk guidelines. The majority of detected policy violations fell into categories such as O3: nudity or O4: sexual content, followed by O2: violence, harm or cruelty, and O6: weapons or substance abuse. Exemplary images are shown in Fig. 5(b). This detailed overview and categorization of potentially unsafe outputs underscore the necessity for robust, context-sensitive safeguards in AI deployments.
@@
Finally, we conducted manual probing on the model safeguarding experiment and largely agree with LlavaGuard’s assessment. Unfortunately, these results further confirm observations of previous works [5, 36, 10] that sexually explicit and nude imagery of women is remarkably easy to produce with seemingly safe prompts. This behavior urges more research into safe generative models and the development of safety guardrails.
@@ 7 Conclusion
We introduced LlavaGuard, a suite of safeguards based on Llava VLMs, designed for assessing images’ safety. LlavaGuard goes beyond rigid classifications and provides assessments that include violated categories and detailed rationales. We also introduced a safety risk taxonomy for assessing the safety of images alongside a human-annotated safety dataset labeled based on this taxonomy. LlavaGuard is tuned on this novel dataset, incorporating varying safety policies. We validated LlavaGuard’s performance on a held-out test set, in which even our smallest model, LlavaGuard-7B, outperforms the much larger Llava-34B baseline. Lastly, we demonstrated LlavaGuard’s efficacy for dataset auditing and model safeguarding. We believe LlavaGuard serves as a strong cornerstone for VLM-based content moderation and beyond. For future work, LlavaGuard would generally benefit from extending its training and test data, specifically with synthetic content. Another promising area for exploration involves extending the categories to encompass bias assessment to promote fairness. Finally, applying LlavaGuard to more potent pre-trained VLMs and auditing other datasets represent intriguing future directions.
Limitations. During the tuning process of LlavaGuard models, human supervision was applied solely to category and rating entries, while the rationales were generated synthetically. Additionally, the annotation of our dataset was largely guided by the default policy outlined in App. A. We encourage future work to explore annotations that consider varying policies. The computational cost is still significant compared to lightweight classifiers, especially for auditing large datasets. Hence, more investigation is needed into improving LlavaGuard’s efficiency, e.g. applying quantization. For the safety taxonomy, we acknowledge the limited presentation in this work. As highlighted previously, safety is highly context- and situation-dependent, making a single general definition often impractical. However, we argue that, much like in law-making, defining a set of general rules for safety is reasonable. Furthermore, our LlavaGuard models are currently all based on Llava. Yet, they are agnostic to the underlying VLM, and newer models can be integrated easily.
Societal Impact. LlavaGuard generally promotes safety for visual datasets and generative models. However, as with any tool, it may also face dual use. First, it might be misused to intentionally obtain unsafe content only, instead of filtering it out. Furthermore, it might be misused to do adversarial content moderation, e.g. suppress content from marginalized groups or ban certain topics (oppressing freedom of speech or press). Another trade-off needing consideration is determining the threshold between safe and unsafe. The choice of this threshold depends on the specific use case, whether prioritizing higher recall or specificity (cf. App. Tab. 4). Future work should explore this threshold in more detail.
@@ Appendix B Training LlavaGuard
We tune all models for a total of 3 epochs on our augmented training set employing a learning rate of $2\mathrm{e}{-5}$, using a cosine scheduler with a warm-up phase of $0.05\%$ steps. We adjusted batch sizes according to the models’ scale: 12, 3, and 8 samples per device for the 7B, 13B, and 34B models, respectively. For 34B, we employed CPU offloading to manage resource requirements. Individual training runs were executed on 4 A100-SXM4-80GB GPUs, each taking less than 4 hours to complete.
Appendix C More qualitative experiments
(a) Qualitative evaluation of the Llava-13B
(b) Qualitative evaluations of LlavaGuard-13B

### v2

@@ 6 Experimental Evaluation
We start with a comprehensive evaluation of LlavaGuard. First, we present qualitative examples to understand potential use cases and the quality of LlavaGuard’s safety assessments. Next, we explore the inferior performance of state-of-the-art safeguards and VLMs compared to LlavaGuard, before turning to one of LlavaGuard’s practical applications in the following section.
Qualitative Results. Fig. 2 presents qualitative examples from the LlavaGuard test set. LlavaGuard’s assessments include safety rating, category, and rationale. Our model not only assigns accurate ratings and categories but also demonstrates transparent policy-following capabilities within the rationales. Specifically, LlavaGuard utilizes the defined policies to assess images, clearly explaining how and why each image complies with or violates the risk guidelines. Furthermore, when the safety policy is modified, LlavaGuard appropriately adjusts its assessments—changing the safety rating from unsafe to safe—and provides solid justifications for these changes in the rationale. In App. Fig. 7, we extend our qualitative evaluation by comparing the assessments of LlavaGuard and corresponding Llava base model using additional images from our test set. We observe that while the base model effectively identifies the content of the images, it fails to adhere to the safety policy. In fact, the policy appears to have no major impact on the base model’s evaluations; it does not account for the policy guidelines within its rationale, nor is it able to adjust its assessment when the policy is modified. In contrast, LlavaGuard provides consistent assessments across these examples and continues to demonstrate strong policy-following capabilities, providing well-grounded reasoning using the risk guidelines of the relevant safety category. Additionally, it demonstrates excellent responsiveness to policy changes.
@@
Overall, our analyses stress distinct features of LlavaGuard: its open-ended rationale generation, which enhances assessment understanding and transparency, and its commonsense capability, which facilitates flexible policy adjustments.
Table 2: Performance comparison of LlavaGuard with Llava-OV (Li et al., 2024c), GPT-4o, OpenAI Omni-Mod (OpenAI, 2024b), LlamaGuard-3-11B (Chi et al., 2024), and ImageGuard (Li et al., 2025) on the held-out test set. We report balanced accuracy, recall, precision, and policy exception score. LlavaGuard substantially outperforms its baselines and SOTA models. Best values are bold, while runner-up is underlined.
Models Open Acc Recall Preci- PES
(%$\uparrow$ (%$\uparrow$ sion (%$\uparrow$ (%$\uparrow$
VLMs Llava-OV-0.5B 52.00 04.23 90.00 68.07
Llava-OV-7B 60.81 29.17 75.00 66.03
GPT-4o222experiments performed with gpt-4o-2024-11-20 72.92 55.99 81.05 77.29
Safeguards LlamaGuard-3-11B 50.28 00.56 100.0 66.92
OpenAI-omni-mod. 66.92 45.24 47.50 60.23
ImageGuard 70.98 83.33 60.98 27.00
LlavaGuard-0.5B 88.70 86.67 87.89 87.10
LlavaGuard-7B 90.84 91.39 87.97 89.85
Empirical Results. In Tab. 2, we expand upon previous qualitative findings and compare LlavaGuard’s performance on our held-out test set to its baseline VLMs and state-of-the-art safeguards. First, both LlavaGuard models consistently outperform their baselines, improving balanced accuracy by more than 30% compared to Llava-OV (Li et al., 2024c). Furthermore, while multiple other SOTA safeguards demonstrate basic safety understanding in images, their overall ability to evaluate safety is strongly limited. For example, the Meta’s moderation tool Llama-Guard-3-11B-Vision (Chi et al., 2024), has an almost negligible recall, misclassifying nearly all images as ‘safe,’ rendering it unreliable for assessing image safety. Even more strikingly, despite being explicitly designed for this task, both ImageGuard (Li et al., 2025) and OpenAI’s omni moderation (OpenAI, 2024b) achieve only around 70% accuracy–significantly underperforming compared to LlavaGuard. Additionally, LlavaGuard is the only model that demonstrates the desired ability to discern and reject unsafe visual content as evidenced by its recall performance along with high accuracy.
Second, LlavaGuard is able to effectively adjust its safety assessment to various policies, as evident by the policy exception score (PES), cf. Eq. 2. Even our smallest model, LlavaGuard-0.5B, outperforms all other VLMs and safeguards. In stark contrast to LlavaGuard, ImageGuard fails to adapt to different policy specifications, as reflected by its low PES (cf. Tab. 2). This limitation persists even when evaluating ImageGuard’s default policy (cf. App. Tab. 6). Overall, ImageGuard shows strong overfitting to one fixed policy, raising substantial concerns about its practical applicability in real-world, dynamic regulatory environments.
In Fig. 3, we evaluate the performance of LlavaGuard and selected VLMs and safeguards across individual safety categories. Consistent with previous findings, LlavaGuard maintains superior performance across categories. In contrast, all other models exhibit substantial inconsistencies in their performance across categories. For example, ImageGuard demonstrates particularly weak recall in category O1.
(Un)ambiguous Cases. Having established LlavaGuard as the top-performing model, we now dive deeper into its performance. Specifically, we evaluate its performance on edge cases (ambiguous) near the unsafe/safe boundary, as well as on unambiguous cases that are clearly distant from this boundary. While it is expected that the performance improves slightly with more distinct cases, the gap is minimal (cf. App. Tab. 4), showing that LlavaGuard handles even edge cases effectively. This observation is essential, demonstrating that LlavaGuard models have successfully captured key characteristics of image safety, making them well-suited for real-world applications, as discussed next.
Figure 3: Category-wise analysis of safety performance. LlavaGuard shows consistent coverage of safety categories whereas other models exhibit either overall or category-specific limitations.
@@
Refer to captionCloakO1 Refer to captionTuskerO8 Refer to captionCleaverO2
Refer to captionBrassiereO3 Refer to captionTubO4 Refer to captionTubO2
@@
7 LlavaGuard: Applied Use Cases
Following up on the general performance evaluation, we now look into two key, real-world use cases of LlavaGuard: (i) dataset auditing and (ii) safeguarding generative models.
Dataset Auditing. In the context of dataset auditing (Gebru et al., 2021), LlavaGuard serves as an annotation tool to identify and document risks associated with unsafe content in large-scale dataset. This helps ensure the integrity and safety of data and downstream AI models (Schramowski et al., 2022; Birhane et al., 2021).
To demonstrate this application, we start by auditing the ImageNet (cf. Fig. 4) dataset with LlavaGuard. Further datasets documentations, namely CC12M (Changpinyo et al., 2021a), COCO (Lin et al., 2014a) and Stylebreeder (Zheng et al., 2024), can be found in App. C.
Fig. 4 illustrates that LlavaGuard assigns 105k images out of 1.3M from ImageNet to one of the 9 safety categories. Among these, 20k instances (19% of the subset and 1.5% of the entire ImageNet) violate the safety policy, thus are rated as unsafe. On the other hand, the vast majority of ImageNet (98.5%) adheres to the safety standards and was rated as safe. Many images fall under category O6: Weapons or Substance Abuse which is a result of ImageNet classes ‘assault_rifle’, ‘tank’, and ‘rifle’. Yet, LlavaGuard clearly distinguishes between guideline-violating images (only 11k out of 36k). These findings, are in line with previous works (Schramowski et al., 2022; Birhane & Prabhu, 2021) which have also identified a substantial number of potentially unsafe images in ImageNet. For example, Schramowski et al.’s classifier flagged over 40k unsafe images. However, upon manual inspection, we found their classifier to be more conservative, failing to differentiate between benign depictions of weapons and illegal ones.
@@ 7 LlavaGuard: Applied Use Cases
Refer to captionO6 Refer to captionO2 Refer to captionO4
Refer to captionO3 Refer to captionO5 Refer to captionO1
@@
Figure 5: Safeguarding generative models. LlavaGuard applied to I2P (11k images generated with StableDiffusion-v1.5). In summary, LlavaGuard successfully detects synthetic candidate images and categorizes them as un/safe according to its taxonomy. (a) reports quantitative results encompassing overall category detections as well as the portion classified as unsafe. The results are also split by category. LlavaGuard performs well in the safety assessment of synthetic content. (b) illustrates examples of images classified as unsafe, with the safety category shown in red.
In Fig. 4(b), we present examples of unsafe images from ImageNet. These samples are clearly unsafe and violate the safety policy. In addition, the assigned safety categories are well aligned with the depicted content. These examples underscore a critical challenge with large-scale datasets: while the general use of these images may be problematic, the human-assigned labels are often even more questionable. In more detail, the labels assigned often do not align with the core content of the image. For instance, the image at bottom center is labeled as ‘bath tub’, yet it primarily displays explicit nude content (O4). Associations like these can lead to spurious, harmful correlations in models trained on this data. These findings underline the need for advanced auditing tools like LlavaGuard in data curation and preprocessing pipelines, especially when handling data at scale where full manual annotation is not feasible.
Model Safeguarding. While dataset auditing can lead to safer models, implementing adequate safeguards during deployment remains crucial. Consequently, we considered StableDiffusion-v1.5 (SD1.5) (Rombach et al., 2022), a model known for its susceptibility to generating unsafe material (Schramowski et al., 2023). We leverage the distilled inappropriate image prompts (I2P) benchmark (Brack et al., 2023b) to elicit the generation of potentially problematic material and subsequently analyze the generated images with LlavaGuard. These prompts (1.1k in total) are specifically designed to evade classical input filters to result in unsafe images. We generated 10 images for each prompt, resulting in 11k images.
The analysis of these generated images (cf. Fig. 5(a)) reveals numerous safety violations (20%). Considering that T2I models are trained on large-scale datasets containing substantial amounts of unsafe content, as observed above, they consequently are able to generate unsafe content across all categories. Especially in the case of nudity (O3), nearly all categorized images (around 90%) also violate the safety policy, indicating the T2I’s model inclination to generate explicit nudity (see left bottom and top right in Fig. 5(b)). Further exemplary images are shown in Fig. 5(b).
@@
To validate LlavaGuard’s assessments, we manually probed the generated images and largely agree. Thus, confirming observations of previous works (Birhane et al., 2021; Schramowski et al., 2023; Brack et al., 2023b) that sexually explicit and nude imagery of women is remarkably easy to produce with seemingly safe prompts. This behavior urges more research into safe generative models and the development of safety guardrails.
Table 3: Agreement rate of human users with LlavaGuard assessments. The high agreement with human assessments shows that LlavaGuard accurately captures safety in vision content across all three dimensions.
human agreement rating category rationale
88% 87% 81%
Measuring Human Agreement with LlavaGuard. To extend previous evidence, we asked human users to annotate LlavaGuard assessments from a broad array of applied use cases. The assessments of LlavaGuard are sampled across a diverse range of datasets—including both real (CC12M (Changpinyo et al., 2021a), COCO (Lin et al., 2014a), ImageNet (Deng et al., 2009)) and synthetic (I2P-generated images (Brack et al., 2023b) by StableDiffusion 1.5, Stylebreeder (Zheng et al., 2024)) datasets—, providing a strong coverage across domains. For further details, see App. J. As shown in Tab. 3, LlavaGuard demonstrates a high level of agreement with human assessments, particularly in safety ratings and category classifications. The agreement is naturally slightly lower for the more complex rationales. These results underscore LlavaGuard’s effectiveness in delivering high-quality, human-aligned evaluations across various applications.
@@ 8 Conclusion
We introduced a novel framework for vision safeguard models, including a safety risk taxonomy for assessing the safety of images alongside a human-annotated safety dataset labeled based on this taxonomy. LlavaGuard goes beyond rigid classifications and provides assessments that include violated categories and detailed rationales. Our empirical results show that LlavaGuard serves as a strong cornerstone for VLM-based safeguarding vision datasets and models. For future work, LlavaGuard would generally benefit from extending its training and test data, specifically with synthetic content. Another promising area for exploration involves extending the categories to encompass bias assessment to promote fairness.
Limitations. During the tuning process of LlavaGuard models, human supervision was applied solely to category and rating entries, while the rationales were generated synthetically. Additionally, the annotation of our dataset was largely guided by the default policy outlined in App. A. While we incorporated policy permutations during training to accommodate diverse policy specifications, we encourage future work to explore annotations that consider varying policies. The tradeoff between computational cost and performance is an important consideration, especially when auditing large-scale datasets and run-time monitoring generative models. To address this, we provide both smaller (0.5B) and larger (7B) checkpoints to accommodate varying requirements. We recognize that this work’s safety taxonomy provides foundational coverage of the safety categories, offering opportunities for further expansion and refinement. As previously mentioned, safety is highly context- and situation-dependent, which makes a single, universal definition and taxonomy often impractical. Yet, we argue that, much like in law-making, our taxonomy adopts a reasonable approach by defining a set of general safety rules. Moreover, LlavaGuard already demonstrates remarkable capabilities in handling different policies, as evidenced by its high PES values. Moreover, as more powerful open VLMs become available, it is expected that LlavaGuard will continue to improve, given that it is agnostic to its underlying VLM.
Impact Statement
LlavaGuard generally promotes safety for visual datasets and generative models. However, as with any tool, it may also face dual use. First, it might be misused to intentionally obtain unsafe content only, instead of filtering it out. While this is helpful for safety research, malicious downstream applications remain. Furthermore, it might be misused to do adversarial content moderation, e.g. suppress content from marginalized groups or ban certain topics (oppressing freedom of speech or press). Another trade-off needing consideration is determining the threshold between safe and unsafe. The choice of this threshold depends on the specific use case, e.g. whether prioritizing higher recall or specificity. Future work should explore this threshold in more detail.
@@ Appendix B Training LlavaGuard
We tune all models for a total of 3 epochs on our augmented training set employing a learning rate of $2\mathrm{e}{-5}$, using a cosine scheduler with a warm-up phase of $0.05\%$ steps. We used a train batch size of 1 with 25 steps of gradient accumulation for both models, 0.5B and 7B, respectively. Individual training runs were executed on 5 A100-SXM4-80GB GPUs, each taking less than 4 hours to complete.
Appendix C LlavaGuard: Further Dataset Audits
Fig. 6 shows further dataset documentations using LlavaGuard.
(a) CC12M
(b) COCO
(c) Stylebreeder
Figure 6: Dataset audit statistics. Next to Imagenet, we documented potentially unsafe content of the vision datasets CC12M (Changpinyo et al., 2021a) and COCO (Lin et al., 2014a) as well as the generative content dataset Stylebreeder (Zheng et al., 2024) using LlavaGuard.
Appendix D Qualitative Comparison: Llava vs. LlavaGuard
(a) Qualitative evaluation of the Llava-ov-7B
(b) Qualitative evaluations of LlavaGuard-7B